# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ A.5 Modified composite law
For $A\leq B$, $\Omega$ does not increase during healing under a constant $\tau$ (eq. A.5). Hence, if the initial condition meets $\Omega<\beta$ as assumed throughout this appendix section, $\Omega$ remains below $\beta$ forever, so that the modified composite law remains to approximate the aging law over the entire time range:
@@
Therefore, the modified-composite-law predictions in constant-$\tau$ healing tests agree with the canonical $B\ln t$ healing for $A\leq B$, as confirmed with the numerical example Fig. A.1(d2&d3). Note that eq. (A.51) is only an order estimate because the condition $\Omega<\beta$ includes the neighborhood of $\Omega=\beta$ as well as $\Omega\ll\beta$, thus only guaranteeing that $f_{t}=\exp(-\Omega_{\rm init}/\beta)>1/e$, whereas the aging law holds that $f_{t}=1$.
For $A>B$, the modified composite law approximates the aging law in the early stage where $\Omega<\beta$ and the slip law in the later stage where $\Omega\gg\beta$. As shown in §A.4, the composite law can saturate before it switches from the slip law to the aging law. In contrast, the modified composite law, which approximates the aging law only for $\Omega<\beta\ll 1$, predicts that saturation ($\Omega$ approaches 1 from below) is necessarily preceded by the switch from aging-law-like to slip-law-like evolution (Fig. A.1, d1). The switch time $t_{\rm as}$ when $\Omega$ reaches $\beta$ is estimated as follows by using eqs. (A.5), (A.12), (A.13), and (A.51):
@@
$\theta_{\rm mod.composite}(t)\sim\theta_{\rm slip}(t-t_{\rm as})|_{\theta(0)=\theta_{\rm aging}(t_{\rm as}),\Omega(0)=\beta}\hskip 10.0pt(A>B\cap t\gg t_{\rm as})$ (A.54)
where we use the same substitution notation as in eq. (A.47). Because evolution after $t_{\rm as}$ is approximated by the slip law, the saturation time $t_{\rm*|mod.composite}$ of $\theta_{\rm mod.composite}$ can be estimated from eq. (A.33) with the initial-value replacement and time translation as per eq. (A.54), considering the saturation time $t_{*}^{\prime}$ for $\theta_{\rm slip}$ is given by eq. (A.33) when the initial condition is $(\theta(0),\Omega(0))=(\theta_{\rm init},\Omega_{\rm init})$:
@@
where we use $\theta_{\rm aging}(t_{\rm as})=t_{\rm as}+t_{\rm c|aging}$ from eq. (A.13).
As long as eqs. (70), (80), and (83) hold, the evolution laws are described by the same asymptotics [$\dot{\theta}\sim 1-\Omega\ln\Omega$ ($\Omega\ll\beta$) and $\dot{\theta}\sim-\Omega\ln\Omega$ ($\Omega\gg\beta$)] as the modified composite law. Therefore, all the approximate behaviors expected of the modified composite law for the constant-$\tau$ hold test, explained in this §A.5, should apply to any evolution laws satisfying eqs. (70), (80), and (83), which define an eligible class of evolution laws (§4.2).
@@ Appendix B Predictions of the slip law during stationary-loadpoint SHS tests
In the present paper, we have taken up SHS||||NSC tests, where slip during the hold phase is too small for the slip law to cause significant healing, as the case which clearly contradicts the prediction by the slip law. In the main text, we did not consider stationary-loadpoint SHS tests at all. Although we emphasize that no existing evolution laws cannot reproduce stationary-loadpoint SHS tests with high machine stiffness (Bhattacharya et al., 2017), it is well known that loadpoint-SHS tests under typical (i.e., not so high) stiffness $k$ of usual apparatus are well reproduced by the slip law (Bhattacharya et al., 2017). As mentioned in §3.3, we did not consider stationary-loadpoint SHS tests as a constraint in deriving the presently proposed evolution law because requirements from the other types of experiments have already constrained the evolution law uniquely. On the other hand, it is necessary to check if our proposed evolution law encompasses the well-established observation that the slip law reproduces healing in stationary-loadpoint SHS tests with typical $k$.
We here examine how the constraints on the evolution law would be modified if stationary-loadpoint SHS tests with typical $k$ are taken to be as an additional requirement. Specifically, we examine how such experiments constrain possible parameter space for the aging-slip switch.
Our conclusions can be grasped from Fig. B.1, which shows the evolution of slip, slip rate, strength, and $\Omega$ against time $t$ during the hold phase of stationary-loadpoint SHS tests. Three cases with $A>B$, $A=B$, and $A<B$ are shown. We assumed $k\sim A/D_{\rm c}$, which is typical in existing experiments. Although $V$ keeps decreasing as $1/t$, $\Omega$ does not decay much; $\Omega$ remains above $10^{-2}$ for $t<1000(V_{\rm prior}/D_{\rm c})$. Because in most stationary-loadpoint SHS tests, $V_{\rm prior}\sim 1$ $\mu$m/s, $D_{\rm c}\sim 1$ $\mu$m, and hold time is up to $1000$ s, this result means that $\Omega$ has been probed only above 0.01. To achieve $\Omega=10^{-2}$ for typical values $A\simeq B$, $t\sim 4\times 10^{6}(V_{\rm prior}/D_{\rm c})$ is necessary. Hence, the range of $\Omega$ probed by stationary-loadpoint SHS tests with typical $k$ ($kD_{\rm c}/\sigma$) and $t$ ($V_{\rm prior}t/D_{\rm c}$) is a mere subset of the range probed in VS tests (discussed in §4.2.1). Therefore, requirement from stationary-loadpoint SHS tests with typical $k$ is automatically satisfied by satisfying the requirement from VS tests. In other words, the slip law’s success in reproducing stationary-loadpoint SHS tests with typical $k$ means nothing more than the slip law’s success in reproducing VS tests.
@@
In the main text, we have pointed out the necessity of aging-slip switch across the border between the regime covered by existing VS tests and that covered by constant-$\tau$-hold SHS tests. The above argument based on Fig. B.1 additionally shows that this aging-slip switch occurs across the border between the regime covered by existing stationary-loadpoint SHS tests with typical $k$ and that covered by constant-$\tau$-hold SHS tests.
In addition, Fig. B.1. shows that $V$ decreases to $V_{\rm prior}/1000$ in $t\sim 1000(V_{\rm prior}/D_{\rm c})$, a realistic experimental time. Hence, typical stationary-loadpoint SHS tests can reject the aging-slip switch at $V_{\rm c}\sim 10^{-8}$ m/s suggested by Kato & Tullis (2001). On the other hand, because of the limited range of reachable $\Omega$ (Fig. B.1), typical stationary-loadpoint SHS tests are difficult to access our proposed condition for the aging-slip switch at $\Omega\sim\beta\lesssim 10^{-2}$.
@@
Initial conditions are $V=V_{\rm prior}$ and $\Omega=1$ [i.e., $\Phi=\Phi_{\rm SS}(V_{\rm prior})$]. Note that we assume $D_{\rm c}=L$ as presumed in the slip law.
In the below, we first solve the coupled differential equations (B.56)–(B.58) and obtain the solutions as functions of $\delta$B.1). Then, in §B.2, we will obtain approximate expressions of those solutions as functions of time, for $\delta\lesssim D_{\rm c}/|1-B/A|$, which holds very well for typical frictional properties $A\simeq B$. §B.3 will discuss behaviors for $\delta\gg D_{\rm c}/|1-B/A|$. However, as will be shown in §B.2, such large $\delta$ is unlikely to be involved in existing stationary-loadpoint SHS tests, and hence their reproducibility is not a concern of the present paper, which aims at reproduction of consensual canons based on large body of experiments. Nevertheless, we include §B.3 to point out that the slip law predicts $A/B$ dependence of $B_{\rm heal}$ in the stationary-loadpoint SHS tests (Fig. B.1), which are partly visible even in the $\delta\lesssim D_{\rm c}/|1-B/A|$ regime. The other aspects of §B.3 have been already discussed by Bhattacharya et al. (2017), but included here for completeness of derivation.
@@ B.1 Exact solutions as functions of slip $\delta$
Equation (B.57) is a linear equation in terms of $\ln V$ and $\Phi$ ($\ln\theta$), and eqs. (B.56) and (B.58) are linear equations in terms of $\tau$, $\ln V$, $\Phi$. Therefore, it is easy to obtain the solutions of the coupled differential equations (B.56)–(B.58) as functions of $\delta$. In the following, we use $\ln\Omega$ instead of $\Phi$ to simplify calculations.
@@
Thus, the slip-law prediction during the stationary-loadpoint hold has been obtained in closed-form functions of slip $\delta$. However, in time $t$ domain, those exact solutions are not expressed in closed forms (exemplified by eq. B.66 for $A=B$ using a special function). Thus, in the following, we will derive approximate closed-form solutions (§B.2 for moderate slip; §B.3 for much large slip).
@@
The lowest-order approximations of (B.62) and (B.63) with respect to $\delta|1-B/A|/D_{\rm c}$ are
@@
$\left(\frac{B\delta}{AD_{\rm c}}+1\right)\simeq\sqrt{1+\frac{2B}{kD_{\rm c}}\left[\ln\left(\frac{t}{t_{\rm c,lph}}+1\right)+\ln\sqrt{1+\frac{2B}{kD_{\rm c}}\ln\left(\frac{t}{t_{\rm c,lph}}+1\right)}\right]+...}$ (B.69)
@@
The first term on the right-hand side of eq. (B.72) represents the lowest-order correction. The right-hand side of eq. (B.71), neglecting the ... term in $\lambda(t)$, serves as an approximate closed-form solution of $\delta(t)$ for $\delta\lesssim D_{\rm c}/|1-B/A|$, the accuracy of which are clear in Fig. B.1 for $\delta\ll D_{\rm c}/|1-B/A|$.
@@
The closed-form approximate solutions eqs. (B.71)–(B.75) [with ... term in eq. B.72 neglected for $\lambda(t)$] are plotted in Fig. B.1 for $A=B$. They agree well with the numerical solutions, demonstrating accuracy of approximation for $\delta\ll D_{\rm c}/|1-B/A|$.
Incidentally, notice Fig. B.1a indicates that the slope $B_{\rm heal}$ is below $B$ for $tV_{\rm prior}/D_{\rm c}<10^{10}$, although eq. B.75 suggests $B_{\rm heal}$ converges to $B$ when $t\to\infty$. Therefore, in loadpoint-hold SHS tests with realistic range of hold time, the slip law predicts that, when $A=B$, then $B_{\rm heal}<B$.
@@ B.3.1 $A>B$ cases
Under this condition of nearly constant $\ln\Omega$, strength recovery is known if $\ln V$ is constrained because $d\ln\Omega=d\ln\theta+d\ln V$. By taking the difference of eq. (B.63) from its value at $t=t^{*}_{\rm lph}$, we have
@@
This is a differential equation for $\delta$ as $V=d\delta/dt$, the solution of which is, as shown in §A.3,
@@
Now we can caluculated the strength recovery as $d\ln\theta\sim-d\ln V$ via eqs. (B.80) and (B.82):
@@ B.3.2 $A<B$ cases
We here derive that the slip law with $A<B$ predicts $A\ln t$ healing (Fig. B.1) for stationary-loadpoint SHS tests with a very long duration, just as we saw for constant-$\tau_{\rm hold}$ SHS tests (§A.3). Throughout the remaining of this section §B.3.2, $A<B$ and $t>t^{*}_{\rm lph}$ are assumed. Under this condition, from eq. (B.62),
$\ln\Omega=\mathcal{O}[\exp[(B/A-1)\delta/D_{\rm c}]],$ (B.85)
@@ Appendix C Predictions of the aging law during VS tests with sufficiently high machine stiffness
Consider ideal VS tests where $\delta=tV_{\rm after}$. For such tests, the early stage following a sufficiently large positive $V$ step (i.e., $V_{\rm after}/V_{\rm before}\gg 1$) achieves $\Omega\gg 1$, where the aging law (eq. 4) predicts $\dot{\theta}\simeq-\Omega$. In the early stage following a sufficiently large negative $V$ step ($V_{\rm after}/V_{\rm before}\ll 1$), $\Omega\ll 1$ holds, and eq. (4) becomes $\dot{\theta}\simeq 1$. Then, further using eq. (11), we find
@@
which indicates strong asymmetry.
@@
Appendix D Derivation of $B=B_{\rm heal}$ from other experimental requirements

### v3

@@ A.5 Modified composite law
For $A\leq B$, $\Omega$ does not increase during healing under a constant $\tau$ (eq. A.5). Hence, if the initial condition meets $\Omega<\beta$ as assumed throughout Appendix A, $\Omega$ remains below $\beta$ forever, and the modified composite law remains to approximate the aging law over the entire time range:
@@
Therefore, the modified-composite-law predictions in constant-$\tau$ healing tests agree with the canonical $B\ln t$ healing for $A\leq B$, as seen in the numerical example Fig. A.1 (d2 & d3). Note, however, that eq. (A.51) is only an order estimate because the condition $\Omega<\beta$ includes the neighborhood of $\Omega=\beta$ as well as $\Omega\ll\beta$, thus only guaranteeing that $f_{t}=\exp(-\Omega/\beta)>1/e$, whereas the aging law holds that $f_{t}=1$.
For $A>B$, the modified composite law approximates the aging law in the early stage where $\Omega<\beta$ and the slip law in the later stage where $\Omega\gg\beta$. As shown in §A.4, the composite law can saturate before it switches from the slip law to the aging law. In contrast, the modified composite law, which approximates the aging law only for $\Omega<\beta\ll 1$, predicts that saturation ($\Omega$ approaches unity from below) is necessarily preceded by the switch from aging-law-like to slip-law-like evolution (Fig. A.1, d1). The switch time $t_{\rm as}$ when $\Omega$ reaches $\beta$ is estimated as follows by using eq. (A.5) with eqs. (A.12), (A.13), and (A.51):
@@
$\theta_{\rm mod.composite}(t)\sim\theta_{\rm slip}(t-t_{\rm as})|_{\theta(0)=\theta_{\rm aging}(t_{\rm as}),\Omega(0)=\beta}\hskip 10.0pt(A>B\cap t\gg t_{\rm as}),$ (A.54)
where the same substitution notation as in eq. (A.47) is used. Because evolution after $t_{\rm as}$ is approximated by the slip law, the saturation time $t_{\rm*|mod.composite}$ of $\theta_{\rm mod.composite}$ can be estimated from eq. (A.33) with the initial-value replacement and time translation as per eq. (A.54), considering the saturation time $t_{*}^{\prime}$ for $\theta_{\rm slip}$ is given by eq. (A.33) with the initial condition of $(\theta(0),\Omega(0))=(\theta_{\rm init},\Omega_{\rm init})$:
@@
where we used $\theta_{\rm aging}(t_{\rm as})=t_{\rm as}+t_{\rm c|aging}$ from eq. (A.13).
Lastly, we note that any evolution laws satisfying eqs. (70), (80), and (83) are described by the same asymptotics [$\dot{\theta}\sim 1-\Omega\ln\Omega$ ($\Omega\ll\beta$) and $\dot{\theta}\sim-\Omega\ln\Omega$ ($\Omega\gg\beta$)] as the modified composite law. Therefore, all the above-derived behaviors in §A.5 should apply to any evolution laws satisfying eqs. (70), (80), and (83), which define an eligible class of evolution laws (§4.2).
@@ Appendix B Predictions of the slip law during stationary-loadpoint SHS tests
In the present paper, we have taken up SHS||||NSC tests, where slip during the hold phase is too small for the slip law to cause significant healing, as the experiments requiring a correction of the slip law. On the other hand, in the main text, we did not consider stationary-loadpoint SHS tests at all because observations in the other types of experiments are enough to constrain the evolution law (§3.3). However, although we have to emphasize that no existing evolution laws can reproduce stationary-loadpoint SHS tests with high machine stiffness (Bhattacharya et al., 2017), it is well known that loadpoint-SHS tests under typical (i.e., not so high) stiffness $k$ of usual apparatus are well reproduced by the slip law not involving purely time-dependent healing (Bhattacharya et al., 2017). One puzzling factor in the aging-versus-slip problem is that some healing experiments require purely time-dependent healing (the aging law) while other healing tests are well described by the slip-driven state evolution (the slip law). We here answer it by examining how the constraints on the evolution law would be modified if stationary-loadpoint SHS tests with typical $k$ are taken as an additional requirement.
Our conclusions can be grasped from Fig. B.1, which shows the evolution of slip, slip rate, strength, and $\Omega$ against time $t$ during the hold phase of stationary-loadpoint SHS tests when the slip law holds. Three cases with $A>B$, $A=B$, and $A<B$ are simulated by assuming $k\sim A/D_{\rm c}$, which is typical in existing experiments. Although $V$ keeps decreasing as $1/t$, $\Omega$ does not decay much; $\Omega$ remains above $10^{-2}$ for $t<1000(V_{\rm prior}/D_{\rm c})$. Because in most stationary-loadpoint SHS tests, $V_{\rm prior}\sim 1$ $\mu$m/s, $D_{\rm c}\sim 1$ $\mu$m, and hold time is up to $1000$ s, this result means that $\Omega$ has been probed only above 0.01. To achieve $\Omega=10^{-2}$ for typical values $A\simeq B$, $t\sim 4\times 10^{6}(V_{\rm prior}/D_{\rm c})$ is necessary. Hence, the range of $\Omega$ probed by stationary-loadpoint SHS tests with typical $k$ ($kD_{\rm c}/\sigma$) and $t$ ($V_{\rm prior}t/D_{\rm c}$), where the slip law is shown to hold, is a mere subset of the range probed in VS tests (discussed in §4.2.1). Therefore, requirement from stationary-loadpoint SHS tests with typical $k$ is automatically satisfied by satisfying the requirement from VS tests. In other words, the slip law’s success in reproducing stationary-loadpoint SHS tests with typical $k$ means nothing more than the slip law’s success in reproducing VS tests.
@@
In the main text, we have pointed out the necessity of aging-slip switch across the border between the regime covered by existing VS tests and that covered by constant-$\tau$-hold SHS tests. The above argument based on Fig. B.1 additionally shows that this aging-slip switch occurs across the border between the regime covered by existing stationary-loadpoint SHS tests with typical $k$ and that covered by constant-$\tau$-hold SHS tests with sufficiently low $\tau_{\rm hold}$.
Besides, Fig. B.1b shows that $V$ decreases to $V_{\rm prior}/1000$ in $t\sim 1000(V_{\rm prior}/D_{\rm c})$, a realistic experimental time. Hence, typical stationary-loadpoint SHS tests can reject the aging-slip switch at $V_{\rm c}\sim 10^{-8}$ m/s suggested by Kato & Tullis (2001). On the other hand, because of the limited range of reachable $\Omega$ (Fig. B.1), typical stationary-loadpoint SHS tests are difficult to access our proposed condition for the aging-slip switch at $\Omega\sim\beta\lesssim 10^{-2}$.
@@
Initial conditions are $V=V_{\rm prior}$ and $\Omega=1$ [i.e., $\Phi=\Phi_{\rm SS}(V_{\rm prior})$]. Appendix B assumes $D_{\rm c}=L$ and $\Omega_{\rm SS}=1$ as presumed in the slip law.
In the below, we first solve the coupled differential equations (B.56)–(B.58) and obtain the solutions as functions of $\delta$B.1). Then, in §B.2, we will obtain approximate expressions of those solutions as functions of time, for $\delta\lesssim D_{\rm c}/|1-B/A|$, which holds very well for typical frictional properties $A\simeq B$. §B.3 will discuss behaviors for $\delta\gg D_{\rm c}/|1-B/A|$. However, as will be shown in §B.2, such large $\delta$ is unlikely to be involved in existing stationary-loadpoint SHS tests, and hence their reproducibility is not a concern of the present paper, which aims at reproduction of consensual canons based on large body of experiments. Nevertheless, we include §B.3 to point out that the slip law predicts $A/B$ dependence of $B_{\rm heal}$ in the stationary-loadpoint SHS tests (Fig. B.1), which are partly visible even in the $\delta\lesssim D_{\rm c}/|1-B/A|$ regime. The other aspects of §B.3 have already been discussed by Bhattacharya et al. (2017), but included here for completeness of derivation.
@@ B.1 Exact solutions as functions of slip $\delta$
Equations (B.56)–(B.58) are linear equations in terms of $\tau$, $\ln V$, and $\Phi$ ($\ln\theta$). Therefore, it is easy to obtain the solutions of the coupled differential equations (B.56)–(B.58) as functions of $\delta$. In the following, we use $\ln\Omega$ instead of $\Phi$ to simplify calculations.
@@
As above, the slip-law prediction during the stationary-loadpoint hold has been obtained in closed-form functions of slip $\delta$. However, in time $t$ domain, those exact solutions are not expressed in closed forms (exemplified by eq. B.66 for $A=B$ using a special function). Thus, in the following, we will derive approximate closed-form solutions (§B.2 for moderate slip; §B.3 for much large slip).
@@
The lowest-order approximations of eqs. (B.62) and (B.63) with respect to $\delta|1-B/A|/D_{\rm c}$ are
@@
$\left(\frac{B\delta}{AD_{\rm c}}+1\right)\simeq\sqrt{1+\frac{2B}{kD_{\rm c}}\left[\ln\left(\frac{t}{t_{\rm c,lph}}+1\right)+\ln\sqrt{1+\frac{2B}{kD_{\rm c}}\ln\left(\frac{t}{t_{\rm c,lph}}+1\right)}\right]+...},$ (B.69)
@@
The first term on the right-hand side of eq. (B.72) represents the lowest-order correction. The right-hand side of eq. (B.71), neglecting the ... term in $\lambda(t)$, serves as an approximate closed-form solution of $\delta(t)$ for $\delta\lesssim D_{\rm c}/|1-B/A|$, the accuracy of which is evident in Fig. B.1 for $\delta\ll D_{\rm c}/|1-B/A|$.
@@
The closed-form approximate solutions eqs. (B.71)–(B.75) [with the ... term in eq. (B.72) neglected for $\lambda(t)$] are plotted in Fig. B.1 for $A=B$. They agree well with the numerical solutions, demonstrating accuracy of approximation for $\delta\ll D_{\rm c}/|1-B/A|$.
Incidentally, notice Fig. B.1a indicates that the slope $B_{\rm heal}$ is below $B$ for $tV_{\rm prior}/D_{\rm c}<10^{10}$, although eq. (B.75) suggests $B_{\rm heal}$ converges to $B$ when $t\to\infty$. Therefore, in loadpoint-hold SHS tests with realistic range of hold time, the slip law predicts that, when $A=B$, then $B_{\rm heal}<B$.
@@ B.3.1 $A>B$ cases
Under this condition of nearly constant $\ln\Omega$, strength recovery is known if $\ln V$ is constrained because $d\ln\Omega=d\ln\theta+d\ln V$. By taking the difference of eq. (B.63), where the second term is proportional to $\ln\Omega$ (eq. B.62), from its value at $t=t^{*}_{\rm lph}$, we have
@@
This is a differential equation for $\delta$ using $V=d\delta/dt$, which is solved as follows as in §A.3,
@@
Now we can calculate the strength recovery as $d\ln\theta\sim-d\ln V$ via eqs. (B.80) and (B.82):
@@ B.3.2 $A<B$ cases
We here derive that the slip law with $A<B$ predicts $A\ln t$ healing (Fig. B.1) for stationary-loadpoint SHS tests with a very long duration, just as we saw for constant-$\tau_{\rm hold}$ SHS tests (§A.3). Throughout the remaining of §B.3.2, $A<B$ and $t>t^{*}_{\rm lph}$ are assumed. Under this condition, from eq. (B.62),
$\ln\Omega=\mathcal{O}\{\exp[(B/A-1)\delta/D_{\rm c}]\},$ (B.85)
@@ Appendix C Predictions of the aging law during VS tests with sufficiently high machine stiffness
Consider ideal VS tests where $\delta=tV_{\rm after}$. For such tests, the early stage following a sufficiently large positive $V$ step (i.e., $V_{\rm after}/V_{\rm before}\gg 1$) achieves $\Omega\gg 1$, where the aging law (eq. 4) predicts $\dot{\theta}\simeq-\Omega$. In the early stage following a sufficiently large negative $V$ step ($V_{\rm after}/V_{\rm before}\ll 1$), $\Omega\ll 1$ holds, and eq. (4) becomes $\dot{\theta}\simeq 1$. Then, further using eq. (11), we obtain
@@
which indicates strong asymmetry (Ampuero & Rubin, 2008).
@@
Appendix D Derivation of $B=B_{\rm heal}$ from other empirical formulae